### v2

# MBSS-T1: Model-Based Self-Supervised Motion Correction for Robust Cardiac T1 Mapping

Abstract
T1 mapping is a valuable quantitative MRI technique for diagnosing diffuse myocardial diseases. Traditional methods, relying on breath-hold sequences and echo triggering, face challenges with patient compliance and arrhythmias, limiting their effectiveness. Image registration can enable motion-robust T1 mapping, but inherent intensity differences between time points pose a challenge. We introduce MBSS-T1, a self-supervised model for motion correction in cardiac T1 mapping, constrained by physical and anatomical principles. The physical constraints ensure expected signal decay behavior, while the anatomical constraints maintain realistic deformations. The unique combination of these constraints ensures accurate T1 mapping along the longitudinal relaxation axis. MBSS-T1 outperformed baseline deep-learning-based image registration approaches in a 5-fold experiment on a public dataset of 210 patients (STONE sequence) and an internal dataset of 19 patients (MOLLI sequence). MBSS-T1 excelled in model fitting quality ($R^{2}$: 0.975 vs. 0.941, 0.946), anatomical alignment (Dice score: 0.89 vs. 0.84, 0.88), and expert visual quality assessment for the presence of visible motion artifacts (4.33 vs. 3.38, 3.66). MBSS-T1 has the potential to enable motion-robust T1 mapping for a broader range of patients, overcoming challenges such as arrhythmias and suboptimal compliance, and allowing for free-breathing T1 mapping without requiring large training datasets. Our code will be publicly available upon acceptance.
1 Introduction
Quantitative T1 mapping is an advanced MRI technique designed to precisely measure intrinsic longitudinal relaxation time in myocardial tissue [27]. This method is increasingly recognized as essential for assessing diffuse myocardial diseases such as inflammation, fibrosis, hypertrophy, and infiltration [27, 29]. “Native” T1 mapping, performed without a paramagnetic contrast agent, effectively detects myocardial conditions, including edema, iron overload, infarcts, and scarring [25]. Furthermore, by allowing for the objective quantification of tissue characteristics, T1 mapping enables longitudinal tracking of changes in the myocardium. This feature is precious for clinical trials, where monitoring the efficacy of interventions over time is critical [25].
Figure 1: Schematic description of T1 mapping for a single voxel. (a) T1-weighted myocardial images at 11 sequential time points. (b) Fitting an inversion recovery curve of the longitudinal magnetization $M_{z}$ over different time points $t$, and extracting the corresponding T1 and $M_{0}$ parameters. (c) Displaying the computed T1 map.
Figure 1 presents the process of generating accurate T1 maps. It begins with capturing a series of images at different time points. Each image captures the myocardium at a distinct point during the longitudinal relaxation. These images are then analyzed to calculate T1 relaxation times through signal decay model fitting, resulting in detailed T1 maps that reflect the myocardial tissue properties. However, heart motion, respiration, and spontaneous patient movements can cause significant distortions in T1 maps, affecting their reliability and clinical usefulness and potentially leading to incorrect diagnoses [28]. Echo-triggering is a well-established method for accounting for cardiac motion. Breath-hold sequences, such as the Modified Look-Locker Inversion Recovery (MOLLI) sequence and its variants [24], are often used to minimize respiratory motion artifacts. However, requiring patients to hold their breath presents practical challenges, as not all patients can fully adhere to these instructions. This breath-hold acquisition further limits the number of images that can be captured and makes T1 mapping unsuitable for patients who cannot tolerate breath-holding [24]. Additionally, inadequate echo-triggering due to cardiac arrhythmias can lead to unreliable T1 maps, compromising diagnostic accuracy.
Reducing motion artifacts and enabling free-breathing cardiac T1 mapping can be accomplished through image registration to align scans taken at different time points. However, achieving precise registration is challenging due to inherent complexities in the image data. These complexities include contrast inversion, partial volume effects, and signal nulling for images captured near the zero crossing of the T1 relaxation curve (Fig. 1).
Despite advancements in classical and deep learning-based methods for motion correction and T1 mapping, significant limitations persist. Classical model-based approaches, while effective, are often slow due to their iterative nature [28, 31]. On the other hand, deep-learning-based methods rely heavily on extensively annotated datasets, which are not readily available for all scanners and protocols [8, 32]. Furthermore, these methods typically use either pairwise loss functions based on metrics such as mutual information to guide the registration process [1] or groupwise approaches such as the principal component analysis (PCA), which ignore the signal decay relaxation model [34]. This omission may result in physically unlikely deformations due to changes in image contrast. These limitations hinder the practical applicability and robustness of current methods.
This work presents MBSS-T1, a self-supervised, physically, and anatomically constrained deep learning model for simultaneous motion correction and T1 mapping from either breath-hold or free-breathing acquisitions. Our method integrates physical and anatomical insights to correct motion in T1 mapping. The main innovation lies in incorporating the signal decay model within the network’s architecture in a self-supervised framework while utilizing a pretrained segmentation network to produce segmentations that guide the registration and ensure the anatomical correctness of the deformed images. This approach enables the generation of corrected T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training. This work extends our previous supervised PCMC-T1 model [13], which required a large annotated dataset and extensive training, thereby limiting its generalization capacity across different acquisition settings, such as sequence type and the number of frames acquired.
We have demonstrated the added value of MBSS-T1 in free-breathing cardiac T1 mapping using a 5-fold experimental setup. This was done on a publicly available free-breathing quantitative dataset of 210 patients [4], acquired using the STONE sequence [30], and an in-house dataset of 19 patients acquired using the MOLLI sequence [24], for both free-breathing and breath-hold scans. Our approach was compared to baseline methods for deep-learning-based image registration [3, 15], as well as the T1 maps produced on the scanner. We used the fit quality (r-squared), the Dice coefficient, and the Hausdorff distance as quantitative metrics to assess registration accuracy and clinical usability. We further demonstrated the clinical impact through qualitative expert radiologist assessments.
2 Related Work
Previous works aimed to address the challenge of motion-robust cardiac T1 mapping can be divided into classical approaches and more recent deep-learning-based approaches. Classical approaches have primarily relied on model-based techniques and iterative algorithms to mitigate the effects of intensity variations between the different frames. More recently, deep learning-based methods have emerged, leveraging large datasets and neural network architectures to enhance motion correction and T1 mapping accuracy. In the following, we review these works.
Classical Approaches:
Zhang et al. addressed intensity variations in their pairwise non-rigid elastic registration framework for T1 mapping by employing the normalized gradient fields difference as their similarity metric [33]. Similarly, Roujol et al. presented an adaptive registration method for varying contrast-weighted images. This approach uses a local non-rigid registration framework based on optical flow to simultaneously estimate motion fields and intensity variations, effectively managing significant contrast differences. Additionally, the method incorporates a regularization term to constrain the deformation field through automatic feature tracking [23].
Guyader et al. introduced a groupwise image registration technique for quantitative MRI that utilizes a dissimilarity metric derived from an approximated form of total correlation. This method addresses the challenge of aligning multiple images simultaneously without relying on a reference image, thereby reducing registration bias and improving the consistency of quantitative measurements [10]. Tao et al. introduced a robust motion correction method for myocardial T1 and extracellular volume mapping using a PCA-based groupwise image registration technique, which registers all MOLLI frames simultaneously with a total-correlation-based metric [26]. Similarly, Huizinga et al. developed a PCA-based groupwise image registration technique for quantitative MRI (qMRI), including T1 mapping, to correct misalignment without a reference image, thus avoiding registration bias. Their method registers all images to a mean space by minimizing a PCA-derived cost function, assuming that a low-dimensional signal model can represent intensity changes in qMRI without needing details of the specific acquisition model [16]. Segmentation-based registration techniques address intensity differences between images using anatomical structures to guide the registration. El-Rewaidy et al. utilized a non-rigid active shape model-based framework, registering T1-weighted images and computing residual motion by matching manually annotated myocardium contours. This method improves T1 map quality by addressing motion-induced errors in myocardial segments [4]. Hering et al. further advanced cardiac motion tracking by combining global semantic information from segmentation labels with local distance metrics, improving the alignment of surrounding structures [14].
A key limitation of the methods above is their failure to account for the signal decay model during registration, resulting in potentially non-physical corrected T1 maps. Model-based approaches have been developed to address this, significantly enhancing T1 mapping accuracy. Xue et al. proposed a fast variational non-rigid registration framework using synthetic image estimation to correct motion in myocardial T1 mapping. Motion-free synthetic images, constructed from initial T1 estimates, guide the registration to avoid motion artifacts [31]. Van Heeswijk et al. introduced a model-based alignment of Look-Locker MRI sequences for myocardial scar quantification, using exponential curve fitting errors as the registration metric [6]. Tilborghs et al. applied a robust motion correction method for cardiac T1 and ECV mapping, integrating data-driven initialization with model-based registration within a non-rigid framework. This method iteratively estimates signal decay model parameters, generates synthetic images, and uses these for accurate registration [28]. However, their iterative nature makes these model-based approaches slower than non-iterative methods.
Deep-Learning Approaches:
Deep-learning approaches have also been proposed for motion correction by image registration. These networks typically learn a function that directly outputs a deformation field given an image pair. For instance, Arava et al. cascaded pairwise registration systems recursively for motion correction in T1-weighted images while minimizing the mutual information between each pair of images [1].
Multi-image registration models have been employed by Hanania et al. using a mutual-information-based loss function [12]. Additionally, Gonzales et al. used a multi-scale U-Net to create distance vector fields and applied warping layers to deform feature maps from coarse to fine, which helps to align images with varying intensities by progressively refining the registration and reducing motion artifacts [8, 7]. Li et al. utilized a sparse coding technique to extract the contrast component from the images. Subsequently, they applied a self-supervised deep neural network (DNN) model for image registration [19]. Another approach by Li et al. used spatial-temporal and physical constraints to enforce low-rank and sparsity, with a deep neural network for efficient T1 mapping, denoising, and artifact reduction [18].
Qin et al. used a joint learning of motion estimation and segmentation to improve registration accuracy by leveraging the unsupervised features learned during motion estimation from a large amount of unannotated data [22]. Pan et al. and Hanania et al. employed GANs to generate virtual images after style transfer at different inversion times from a single input image for image registration. The GANs reduce the sensitivity of the models to the inherent intensity differences by standardizing contrast across images, which helps to align images with varying contrasts more accurately [21, 11].
Yang et al. introduced a sequential method for addressing contrast differences in images. This method initially separates intensity changes from different inversion times from the constant anatomical structure. However, it heavily relies on accurately disentangling the anatomical structure from the contrast. Additionally, the registration process only occurs between the isolated anatomical images, disregarding the signal decay model along the inversion time axis. Consequently, this approach could result in physically unrealistic deformations due to the necessity for direct integration with the signal decay model [32].
These deep learning-based methods, however, do not directly consider the signal decay model, which is essential as the image contrast changes according to it. As a result, these methods may produce physically unrealistic deformations. While classical model-based approaches address these issues, their iterative nature makes them relatively slow.
Previously, we introduced PCMC-T1, a method designed explicitly for free-breathing myocardial T1 mapping that incorporates physical constraints into the motion correction process [13]. This deep learning-based approach uses a supervised framework to ensure accurate T1 mapping under free-breathing conditions. However, the practical application of PCMC-T1 is limited by its reliance on large annotated datasets and extensive training, which are not readily available for every scanner and protocol.
This work introduces a self-supervised approach extending our previous PCMC-T1 model. The proposed MBSS-T1 approach enables the generation of corrected T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training.
3 Methods
The T1 relaxation time is described using the Inversion Recovery T1 Relaxation Equation [30]:
$M_{Z}(M_{0},T1,t)=M_{0}\cdot\left(1-2\cdot e^{-t/T1}\right)\,.$ (1)
In this equation, $M_{0}$ represents the equilibrium magnetization, which is the magnetization of the tissue before any relaxation occurs. The parameter $T1$ is the longitudinal relaxation time, indicating how quickly the magnetization returns to its equilibrium state along the longitudinal axis after inversion. The variable $t$ denotes the time at which the magnetization is measured, and $M_{Z}(M_{0},T1,t)$ is the magnetization at time $t$, given the initial magnetization $M_{0}$ and the relaxation time $T1$.
The estimation of the $T1$ and $M_{0}$ parameters is typically performed by solving a least-squares optimization problem [27]. This involves minimizing the sum of the squared differences between the measured magnetization values and the values predicted by the model. The least-squares optimization problem can be formulated as follows:
$\widehat{T1},\widehat{M_{0}}=\mathop{\mathrm{arg\,min}}\nolimits_{T1,M_{0}}\sum_{i=0}^{N-1}\left(M_{Z,i}-M_{0}\cdot\left(1-2\cdot e^{-t_{i}/T1}\right)\right)^{2}$ (2)
where $M_{Z,i}$ represents the observed magnetization values at different time points $t_{i}$, and $N$ is the total number of measurements.
This classical least-squares approach, as well as deep learning methods such as the one described in [9], assumes that the images are motion-free, typically achieved through prior image registration techniques. However, motion can be a significant factor in practice, especially in cardiac imaging. To account for motion, the optimization problem must be extended to include deformation fields, which can be introduced as additional variables in the least-squares problem.
The optimization problem with motion can be formulated as follows:
$\widehat{T1},\widehat{M_{0}},\widehat{\Phi}=\mathop{\mathrm{arg\,min}}\nolimits_{T1,M_{0},\Phi}\sum_{i=0}^{N-1}\left\|M_{Z}(M_{0},T1,t)_{i}-\phi_{i}\circ I_{i}\right\|^{2}\,.$ (3)
In this extended formulation, $\widehat{T1}$ and $\widehat{M_{0}}$ are the estimated relaxation and magnetization parameters, while $\widehat{\Phi}$ represents the estimated deformation fields that account for motion. Specifically, $\phi_{i}$ denotes the $i$-th deformation field, with $\Phi=\{\phi_{i}\}_{i=0}^{N-1}$. $I_{i}$ corresponds to the $i$-th original image, and $\phi_{i}\circ I_{i}$ is the deformed image, reflecting the influence of motion on the observed magnetization values.
Optimizing this equation directly is challenging due to the large number of unknowns linked to the T1 relaxation model and the accompanying deformation fields. This approach can be solved iteratively using classical methods such as those described in previous studies [31, 28], but these methods tend to be slow and prone to getting stuck in local minima.
To overcome these limitations, we reframe the optimization task as estimating the weights of a deep neural network ($\Theta$) that simultaneously predicts the T1 model parameters and spatial transformations from the observed magnetization signals and corresponding time points. This is achieved using two encoder-decoder modules that operate concurrently: one for predicting the T1 parameters (parametric mapping component) and one for predicting the deformation fields (motion component). The objective function to be minimized is expressed as:
$\Theta=\mathop{\mathrm{arg\,min}}\nolimits_{\Theta}\sum_{i=0}^{N-1}\left\|M_{Z}(f_{\Theta}[1]_{M_{0}},f_{\Theta}[1]_{T1},t)_{i}-f_{\Theta}[0]_{i}\circ I_{i}\right\|^{2}\,.$ (4)
In this formulation, $f_{\Theta}[0]$ represents the output of the parametric mapping component, which includes the deformation fields. Specifically, $f_{\Theta}[0]_{i}$ corresponds to the $i$-th deformation field. Meanwhile, $f_{\Theta}[1]$ denotes the output of the motion component, which includes both the T1 map ($f_{\Theta}[1]_{T1}$) and the M0 map ($f_{\Theta}[1]_{M_{0}}$).
3.1 Model Architecture
Figure 2: Overview of the MBSS-T1 model architecture: (a) The motion component is responsible for deformable image registration, predicting deformation fields that align the acquired images across different time points. (b) The parametric mapping component estimates the T1 and M0 parameters and generates synthetic aligned T1-weighted (T1W) images. (c) The segmentation and uncertainty component uses a pre-trained segmentation network to extract the myocardium segmentations, calculates the deformed segmentations, and retains the ones with high confidence. The network aims to minimize the distance between the motion-corrected T1W and synthetic images while maximizing the Dice score between the deformed segmentations.
Figure 2 presents our overall architecture. Our network comprises two U-Net-inspired encoder-decoder modules that operate simultaneously. The network has three main components:
Motion Estimation Component: The upper module, an extension of the VoxelMorph architecture, functions as a multi-image deformable image registration module. It dynamically adjusts the number of registration heads based on the number of recovery inversion time points ($N$). It outputs a specific deformation field $\{\phi_{i}\mid i=0,\ldots,N-1\}$ and registered image $\{R_{i}\mid i=0,\ldots,N-1\}$ for each timestamp $i$. This module leverages the deformation field prediction layers from [2], and the integration layer from [3].
Parametric Mapping Component: The bottom module predicts the parameters of the qMRI signal relaxation model (T1 and $M_{0}$) and forwards these parameters to the signal generation layer. This layer predicts a sequence of synthetic images $\{S_{i}\mid i=0,\ldots,N-1\}$ following the Inversion Recovery T1 Relaxation Equation (Eq. 1). These images are anatomically aligned according to the correct signal decay model, as each image prediction utilizes the same T1 and $M_{0}$ parameters, with the only difference being the timestamp of each image.
Both encoder-decoder modules follow a U-Net-like structure, including convolutional and downsampling layers and skip connections that establish links between each module’s encoder and decoder.
Segmentation and Uncertainty Component: To ensure the anatomical correctness of the aligned images, we added anatomical constraints to the registration process. First, nnU-Net [17] (other segmentation networks can also be used) is utilized as a pretrained network to predict the myocardium in every unregistered frame. Second, an uncertainty block is employed to retain only the segmentations with high confidence in their correctness. This block receives the masks and softmax maps generated by the nnU-Net output. It retains only those masks that meet two criteria: (1) a required percentage $\gamma$ of the myocardial segmentation pixels must have a score greater than a confidence threshold $\alpha$, and (2) the segmentation mask must consist of exactly one connected component. Here, $P_{i}(x)$ denotes the softmax probability at pixel $x$ in the $i$-th segmentation, and $S_{i}$ is the corresponding segmentation mask. Let $G$ be the set of indices of segmentations that satisfy both of these criteria:
$G=\left\{i\mid\frac{1}{|S_{i}|}\sum_{x\in S_{i}}\mathbf{1}\left[P_{i}(x)>\alpha\right]\geq\gamma\text{ and }S_{i}\text{ is connected}\right\}\,.$ (5)
Next, the masks are aligned using the deformation fields $\Phi=\{\phi_{i}\}_{i=0}^{N-1}$ from the Motion Estimation Component to obtain deformed segmentations. During registration, the Dice score is maximized between the deformed myocardium segmentations to ensure that the myocardiums are properly registered. We define a reference image index $r$ as the time point corresponding to the segmentation with the highest confidence score:
$r=\underset{i\in G}{\arg\max}\left(\frac{1}{|S_{i}|}\sum_{x\in S_{i}}P_{i}(x)\right)\,.$ (6)
This equation selects the index $r$ where the average confidence across the segmentation $S_{i}$ is the highest among the segmentations in the set $G$.
3.2 Loss Functions
We encourage predictions of physically and anatomically plausible deformation fields by coupling three terms in our loss function as follows:
$\mathcal{L}_{total}=\lambda_{1}\cdot\mathcal{L}_{fit}+\lambda_{2}\cdot\mathcal{L}_{smooth}+\lambda_{3}\cdot\mathcal{L}_{Seg}$ (7)
The first term ($\mathcal{L}_{fit}$) addresses the differences between the images predicted by the model’s decoder and the deformed images. To achieve this, we compute the mean squared error (MSE) between the registered images ${R_{i}|i=0,\ldots,N-1}$ and the synthetic images ${S_{i}|i=0,\ldots,N-1}$:
$\begin{split}\mathcal{L}_{fit}(T_{1},M_{0},t_{i=0}^{N-1},\Phi)=&\sum_{i=0}^{N-1}(S_{i}-R_{i})^{2}\\ =&\sum_{i=0}^{N-1}\left(M_{0}\cdot(1-2\cdot e^{\frac{-t_{i}}{T_{1}}})-R_{i}\right)^{2}\,.\end{split}$ (8)
Here, $S_{i}$ denotes the images generated using the signal model equation (Eq. 1), while the registered images are obtained from the Parametric Mapping component. This term enforces the creation of deformation fields that are physically plausible by ensuring consistency with the T1 signal relaxation model.
The second term ($\mathcal{L}_{smooth}$) encourages the creation of realistic and smooth deformation fields $\Phi$. This is achieved by penalizing large $l_{2}$ norms of the gradients of the velocity fields [2]:
$\mathcal{L}_{smooth}(\Phi)=\sum_{i=0}^{N-1}\frac{1}{\Omega}\sum_{p\in\Omega}\left\lVert\nabla\phi_{i}(p)\right\rVert^{2}\,.$ (9)
Here, $\Omega$ denotes the domain of the velocity field, and $p$ represents the voxel locations. By minimizing this term, the model promotes smoother deformation fields, reducing unrealistic distortions.
The last term ($\mathcal{L}_{Seg}$), inspired by [2], encourages anatomically correct deformation fields by maximizing the Dice score between the deformed myocardium segmentations and by maximizing the Dice score between the deformed left ventricle segmentations. We denote $\mathcal{L}_{Seg}$ as follows:
$\mathcal{L}_{Seg}=\mathcal{L}_{MyoSeg}+\mathcal{L}_{LVSeg}\,.$ (10)
Our approach focuses on the segmentations retained in the set $G$, which contains only those segmentations with high confidence. The $\mathcal{L}_{MyoSeg}$ and $\mathcal{L}_{LVSeg}$ segmentation loss functions are therefore defined using only these high-confidence segmentations:
$\begin{split}\mathcal{L}_{\text{MyoSeg}}&(r,\{MyoSeg_{i}\}_{i\in G},\{\phi_{i}\}_{i\in G})=\\ &\sum_{\begin{subarray}{c}i\in G\\ i\neq r\end{subarray}}\text{DiceLoss}(MyoSeg_{r},MyoSeg_{i}\circ\phi_{i})\end{split}$ (11)
$\begin{split}\mathcal{L}_{\text{LVSeg}}&(r,\{LVSeg_{i}\}_{i\in G},\{\phi_{i}\}_{i\in G})=\\ &\sum_{\begin{subarray}{c}i\in G\\ i\neq r\end{subarray}}\text{DiceLoss}(LVSeg_{r},LVSeg_{i}\circ\phi_{i})\end{split}$ (12)
where $MyoSeg_{i}$ is the $i$-th binary segmentation mask of the myocardium, $LVSeg_{i}$ is the $i$-th binary segmentation mask of the left ventricle, and $Seg_{r}$ is the binary segmentation mask of the $r$-th image, which is the reference image. Only the segmentations indexed by $G$, which meet the high-confidence criteria, are used to calculate the loss.
3.3 Implementation details
We implemented our models in PyTorch. Experimentally, we zeroed the $r$-th deformation field and predicted deformation fields only for the rest of the time points. We optimized our hyperparameters using a grid search. The final settings for the loss function parameters were: $\lambda_{1}=1$, $\lambda_{2}=5000$, $\lambda_{3}=80000$, $\gamma=0.99$, and $\alpha=0.9$. We used a batch size of 1 and the ADAM optimizer with a learning rate of $2\cdot 10^{-3}$. Our network overfits each case. For each case, we trained the network from scratch for a fixed number of 40 iterations, which took approximately 6 seconds. We performed hyperparameter optimization for the baseline methods using a grid search. All experiments were run on an NVIDIA Tesla V100 GPU with 32GB of RAM.
4 Evaluation Methodology
4.1 Datasets
4.1.1 Free Breathing - STONE Dataset
A publicly available myocardial T1 mapping dataset [4, 5] consists of 210 subjects, 134 males and 76 females, aged $57\pm 14$ years. All subjects were diagnosed with or suspected of having cardiovascular diseases. The imaging was performed on a 1.5T MRI scanner (Philips Achieva) equipped with a 32-channel cardiac coil, using an ECG-triggered, free-breathing, slice-interleaved T1 mapping sequence (STONE) [30]. The acquisition parameters were as follows: field of view (FOV) $=360\times 351$ [mm2] and voxel size $=2.1\times 2.1\times 8$ [mm3]. Five slices were captured from the base to the apex in the short-axis view for each subject at 11 distinct time points. Additionally, the dataset included manual expert segmentations of the myocardium [4]. The images were resized to $160\times 160$ pixels per time point, and min-max normalization was applied to the entire sequence to standardize the image intensities.
4.1.2 Free Breathing and Breath Hold - MOLLI Dataset
Data was acquired on a 3T Siemens MRI system (PRISMA, Siemens Healthineers, Erlangen, Germany) equipped with a 32-channel body coil, using an ECG-triggered, free-breathing, slice-interleaved T1 mapping MOLLI sequence (MyoMaps, SIEMENS Healthcare). The acquisition parameters were as follows: field of view (FOV) $=306\times 360$ [mm2], flip angle of 35°, and an 8 mm gap between slices. Five slices were captured from the base to the apex in the short-axis view for each subject at eight distinct time points. The dataset consists of cardiac images obtained from 19 patients undergoing two scans: one during breath-hold and another during free breathing. Additionally, extra breath-hold and free-breathing scans were performed for five patients for test-retest analysis.
4.2 Experiments
4.2.1 Exp. 1: Motion Correction in Free-Breathing MRI Using STONE Dataset
Experimental setup: Our MBSS-T1 method was compared against two variations: a multi-image registration model using a mutual-information-based loss function (REG-MI) [12], and PCMC-T1, which incorporates physical constraints into the motion correction process within an unsupervised framework [13]. Additionally, two state-of-the-art deep learning algorithms for medical image registration were included in the comparison: pairwise probabilistic diffeomorphic VoxelMorph with a mutual-information-based loss [3] and pairwise SynthMorph [15]. For the self-supervised approaches, MBSS-T1 and SynthMorph, the entire STONE dataset was utilized for inference since no training phase was required. For the unsupervised approaches (PCMC-T1, REG-MI, VoxelMorph) and the nnUNet pretrained segmentation network (used for MBSS-T1), a 5-fold cross-validation strategy was implemented. In each fold, the 210 subjects were divided into 80% for training and 20% for testing. T1 maps generated directly from the acquired images were also used for comparison by applying a standard T1 mapping procedure. A T1 relaxation curve was fitted to the MRI data using the conventional inversion recovery model without any pre-motion correction techniques. Fitting was performed per pixel across the series of T1-weighted images, resulting in maps representing the T1 relaxation times. These maps, referred to as ’w.o motion correction,’ serve as a baseline to evaluate the impact of motion correction on the accuracy and quality of the T1 mapping. The T1 maps generated by our MBSS-T1 model were quantitatively evaluated against those produced both before (’w.o motion correction’) and after deep-learning-based image registration as a pre-processing step. The evaluation metrics used were the $R^{2}$ of the model fit to the observed data in the myocardium, the Dice score, and the Hausdorff distance values of the myocardium segmentations.
Clinical impact: The clinical implications of our method were assessed through a semi-quantitative ranking of the T1 maps for motion artifacts. This assessment was conducted by an expert cardiac radiologist with five years of experience (I.V), blinded to the methods used to generate the maps. A random selection of 29 cases (each with five slices) from the test set and their corresponding T1 maps was evaluated. The radiologist rated each slice, assigning a score of 1 to slices without significant visible motion artifacts and a score of 0 to those with artifacts. Overall patient scores were calculated by summing the scores of the slices. The highest possible score per subject was 5, indicating no motion artifacts in all slices, while the lowest was 0, indicating artifacts in all slices. Statistical significance was evaluated using the repeated measures ANOVA test, with p$<$0.05 considered significant.
4.2.2 Exp. 2: Motion Correction in Free-Breathing and Breath-Hold MRI Using MOLLI
Experimental Setup: The objective of this experiment was to evaluate the effectiveness of the MBSS-T1 method in accounting for motion during free-breathing and breath-hold MRI scans using the MOLLI sequence. We utilized the pre-trained nnUNet model trained on the STONE dataset, as described in the previous section. To make the MBSS-T1 model compatible with the MOLLI sequence, we modified the MBSS-T1 model by adjusting the number of heads in the network output from 11 to 8. The performance of three motion-correction methods was compared: (1) without motion correction, (2) Siemens MyoMaps, and (3) MBSS-T1. For Siemens MyoMaps, we used the motion-corrected T1 maps produced by the scanner’s image processing software.
Clinical Impact: As in Exp 1, the clinical impact of our method was assessed through a semi-quantitative ranking of the T1 maps for the presence of motion artifacts. This assessment was conducted on free-breathing and breath-hold scans to determine how effectively each method mitigated motion artifacts. The primary evaluation metric was a clinical score assigned by an experienced cardiac radiologist with five years of experience based on the clarity and quality of the T1 maps produced by each method.
4.2.3 Exp. 3: Consistency of T1 Mapping with Motion Correction in Test/ReTest MOLLI Dataset
Experimental Setup and Objective: This experiment aimed to assess the consistency and reproducibility of T1 mapping when motion correction is applied. This analysis aimed to evaluate whether motion correction techniques improve the reliability of myocardial T1 measurements across repeated scans, particularly in the presence of motion. The subset included data from five patients undergoing two free-breathing and two breath-hold scans. The goal was to compare the consistency of T1 measurements across these scans using different motion correction methods.
Quantitative Evaluation: To quantitatively assess the reproducibility of T1 measurements, we calculated the intraclass correlation coefficient (ICC3) [20] for mean myocardial T1 values across three different methods: without motion correction, Siemens MyoMaps, and MBSS-T1. The analysis was conducted using data from five patients in the MOLLI dataset, each of whom underwent two free-breathing and two breath-hold scans. For each scan and method, mean myocardial T1 values were computed for four regions of interest (ROIs) – upper, right, lower, and left – on the mid-slice of the T1 map. The ICC3 metric was then calculated for all paired test-retest ROIs to assess the agreement between repeated measurements. A high ICC3 value indicates strong agreement between the test and retest measurements, suggesting robustness and reproducibility in quantifying myocardial T1 values. This metric provides valuable insights into the reliability and consistency of T1 measurements across different methods and scanning conditions. A comprehensive understanding of each method’s performance in clinical practice can be achieved by quantifying the agreement between test and retest measurements across different ROIs and scanning conditions. This analysis also enables the evaluation of the effectiveness of motion correction techniques in improving measurement accuracy and reproducibility.
5 Results
5.1 Exp. 1: Motion Correction in Free-Breathing MRI Using STONE Dataset
Quantitative evaluation: Table 1 summarizes our results for the test sets across all folds, encompassing a total of 210 patients. The MBSS-T1 method outperformed other state-of-the-art registration methods in terms of $R^{2}$, Dice score, and Hausdorff distance. Specifically, MBSS-T1 achieved an $R^{2}$ value of $0.975\pm 0.05$, which represents an improvement of approximately 3% compared to the next best method, PCMC-T1, with an $R^{2}$ of $0.955\pm 0.078$. Additionally, MBSS-T1 produced a Dice score of $0.89\pm 0.075$, surpassing SynthMorph’s $0.88\pm 0.149$, and reduced the Hausdorff distance to $6.43\pm 5.54$ mm, which is a 31% improvement compared to SynthMorph’s $8.59\pm 9.98$ mm. These results suggest that our approach enhances the physical plausibility of deformations through signal relaxation and anatomical consistency. Our results are particularly noteworthy, considering we compared methods requiring a full training process.
Clinical impact: Figure 3 presents several representative cases. The maps generated by the MBSS-T1 method demonstrate the highest quality, as evidenced by superior metrics. The rightmost column of Table 1 summarizes the clinical impact assessment results for the MBSS-T1 method, which achieved the highest quality score of $4.33\pm 0.54$. This score reflects an improvement of approximately 10% over the following best method, PCMC-T1, which had a score of $3.93\pm 0.78$. The difference in radiologist grading was statistically significant (p $\ll 10^{-5}$). Furthermore, as demonstrated in Fig. 5, the MBSS-T1 method achieves a closer fit to the expected magnetization curve (with $R^{2}=0.99$) and reduces artifacts, which may contribute to improved diagnostic confidence. These results suggest the MBSS-T1 method’s potential utility in clinical settings where motion artifacts can impact the accuracy of T1 mapping.
5.2 Exp. 2: Motion Correction in Free-Breathing and Breath-Hold MRI Using MOLLI
The results of Exp 2 indicate that the MBSS-T1 method improves the quality of T1 maps by effectively reducing motion artifacts, as shown by the highest clinical scores in both breath-hold (BH) and free-breathing (FB) scans. Specifically, MBSS-T1 achieved mean scores of $4.02\pm 0.485$ for BH and $4.13\pm 0.495$ for FB scans (Table 2). These scores represent an improvement of approximately 5% in BH and 6% in FB scans compared to Siemens Myomap, which had scores of $3.84\pm 0.472$ and $3.92\pm 0.533$, respectively.
The images in Fig. 4 illustrate the reduction in motion-related distortions achieved by the MBSS-T1 method. This improved motion correction results in more precise and reliable T1 maps with fewer artifacts.
5.3 Exp. 3: Consistency of T1 Mapping with Motion Correction in Test/ReTest MOLLI Dataset
The results in Table 3 show the ICC3 metrics for mean and median myocardial T1 values across three different motion correction methods: without motion correction, Siemens Myomap, and MBSS-T1.
For both mean and median myocardial T1 values, the MBSS-T1 method consistently exhibited higher ICC3 values than the other methods. Specifically, for the mean T1 values, the MBSS-T1 method attained ICC3 values of $0.23$ for free-breathing (FB), $0.42$ for breath-hold (BH), and $0.31$ for combined FB and BH scans. In contrast, the w.o motion correction and Siemens Myomap methods displayed significantly lower ICC3 values, suggesting less agreement between test and retest measurements. Similarly, for the median T1 values, the MBSS-T1 method achieved ICC3 values of $0.26$ (FB), $0.38$ (BH), and $0.3$ (FB+BH), surpassing the other methods.
These findings suggest that the MBSS-T1 method offers more reliable and consistent myocardial T1 measurements than the other methods. However, it is important to note that while the MBSS-T1 method shows the highest ICC3 values, the scores indicate only moderate reliability. Although the MBSS-T1 method improves the robustness and reproducibility of T1 mapping under free-breathing and breath-hold conditions, further improvements are needed to achieve high reliability.
Table 1: Quantitative comparison between motion correction methods for STONE myocardial T1 mapping. Results are presented as mean±$\pm$±std.
$R^{2}\uparrow$ $DSC\uparrow$ HD [mm] $\downarrow$ Clinical Score $\uparrow$
w.o motion correction $0.911\pm 0.12$ $0.664\pm 0.23$ $14.93\pm 11.76$ $2.9\pm 0.92$
SynthMorph $0.946\pm 0.09$ $0.88\pm 0.149$ $8.59\pm 9.98$ $3.66\pm 0.83$
Voxelmorph-seg $0.941\pm 0.096$ $0.84\pm 0.188$ $9.39\pm 11.93$ $3.38\pm 0.65$
Reg-MI $0.95\pm 0.08$ $0.73\pm 0.168$ $16.29\pm 11.43$ $3.68\pm 0.83$
PCMC-T1 $0.955\pm 0.078$ $0.835\pm 0.137$ $9.34\pm 7.85$ $3.93\pm 0.78$
MBSS-T1 0.975$\pm$0.05 0.89$\pm$0.075 6.43$\pm$5.54 4.33$\pm$0.54
Figure 3: Representative T1 maps computed with the different approaches. Our approach (MBSS-T1) demonstrates a clearer delineation between the blood and the muscle with a reduced partial volume effect, resulting in a more homogeneous mapping of the myocardium.
Figure 4: Comparison of T1 mapping results for four patients (rows) using three motion correction methods: Original (left), Siemens MyoMap (middle), and MBSS-T1 (right). MBSS-T1 provides superior motion correction, resulting in clearer T1 maps with fewer artifacts and better differentiation between the blood and myocardium.
Figure 5: Comparison of T1 mapping results with and without motion correction. (Top row) Without motion correction, the T1 fit shows significant deviations in magnetization values (circled), leading to artifacts in the resulting T1 map. This is reflected in the lower $R^{2}=0.69$. (Bottom row) Using the MBSS-T1 method, the T1 fit closely follows the expected curve, resulting in a higher $R^{2}=0.99$ and a more accurate T1 map, with reduced artifacts in the corresponding region. This demonstrates the effectiveness of motion correction in improving the accuracy and quality of T1 mapping.
Table 2: Comparison between motion correction methods for MOLLI myocardial T1 mapping. Results are presented as clinical score mean±$\pm$±std.
Method Clinical Score
$Breath-hold$ $Free-breathing$
w.o motion correction $3.89\pm 0.488$ $3.63\pm 0.742$
Siemens Myomap $3.84\pm 0.472$ $3.92\pm 0.533$
MBSS-T1 $\mathbf{4.02\pm 0.485}$ $\mathbf{4.13\pm 0.495}$
Table 3: Comparison of Interclass Correlation Coefficient (ICC3) Metrics for Mean and Median Myocardial T1 Values
Method ICC3 (on Mean) ICC3 (on Median)
FB BH FB+BH FB BH FB+BH
w.o motion correction $-0.04$ $0.31$ $0.22$ $-0.03$ $0.31$ $0.23$
Siemens Myomap $-0.23$ $0.28$ $-0.008$ $-0.20$ $0.32$ $0.03$
MBSS-T1 $\mathbf{0.23}$ $\mathbf{0.42}$ $\mathbf{0.31}$ $\mathbf{0.26}$ $\mathbf{0.38}$ $\mathbf{0.3}$
6 Conclusions
This work presents a self-supervised method for motion-robust cardiac T1 mapping. The method utilizes a self-supervised learning approach combined with a pre-trained segmentation framework, enhanced by incorporating a signal decay model within the network’s architecture and an anatomical constraint. This approach enables the generation of corrected T1 maps across various imaging protocols without requiring protocol-specific data collection or extensive training. The main contribution of this work lies in the novel combination of signal decay and anatomical constraints within the registration process, providing a powerful tool for improving motion correction and expanding the reach of quantitative MRI techniques in clinical practice.
Compared to baseline methods, the method demonstrated quantitative improvements in fit quality, Dice score, and Hausdorff distance. A qualitative evaluation by an expert cardiac radiologist indicated the clinical relevance of the generated T1 maps, showing potential for accurate T1 estimation in patients unable to perform breath-hold during MRI scans. This approach may extend the applicability of quantitative cardiac T1 mapping across a broader patient population, ensuring reliable imaging even under challenging conditions.
In addition to its effectiveness in T1 mapping, the proposed approach holds significant potential for application to other quantitative MRI problems. By integrating signal decay modeling with anatomical constraints, this method offers a robust framework that could be adapted to correct motion and improve accuracy in various quantitative MRI sequences, such as T2 mapping, diffusion MRI, and beyond. Furthermore, the flexible framework could be further enhanced by replacing the backbone with more advanced architectures, such as transformer-based models. This substitution could leverage the global attention mechanisms inherent in transformers, potentially improving the capture of complex anatomical structures and further enhancing the accuracy of the generated maps.
Acknowledgments
This work was supported in part by research grants from the Israel-US Binational Science Foundation, the Israeli Ministry of Science and Technology, the Israel Innovation Authority, and the joint Microsoft Education and the Israel Inter-university Computation Center (IUCC) program.